Perimeter or a rectangle. The perimeter of a rectangular backyard is 6x + 6 yards. If the width is x yards, find a binomial that represents the length.
Answer by checkley71(8403)

2L+2W=PERIMETER OR 2L+2X=6X+6 OR 2L=4X+6 OR L=(4X+6)/2 OR L=2X+3